OpenStudy (anonymous):

12H+N=210 & 9H+2N/3=155 N=NAILS H=HORSESHOES HOW MANY NAILS WERE IN THE JOB THE BLACKSMITH COMPLETED?

OpenStudy (campbell_st):

well substitution looks easiest take the 1st equation and make N the subject N = 210 - 12H then substitute into the 2nd equation.. distribute, collect like terms and solve for N when you have N substitute into either equation to find H hope this helps

OpenStudy (anonymous):

Ok, you have to use the system of equation. We know equation 1=> 12H + N = 210 equation 2 => 9H + 2N/3 = 155 We have two unknowns H and N with 2 equations so, let's find one unknown first Look at eq. 1 it has N alone so, it's easier to find the equation for N in terms of H 12H + N = 210 N = 210 - 12H Now we have a new equation derived from equation 1. Let's use the new equation in Equation 2. 9H + 2N/3 = 155 9H + (2/3)(210 - 12H) = 155 9H + 140 - 8H = 155 H = 155 - 140 H = 15 Now, we have H and we can find N N = 210 -12H = 210 - 12(15) = 30. Let's check our answers using equation 1. N = 30 , H = 15 12H + N = 210 12(15) + 30 = 210 210 = 210

OpenStudy (anonymous):

thank you =)

OpenStudy (campbell_st):

lol... someone always gives the answer...